OpenStudy (anonymous):

the product of 2 consecutive odd integers is 323 what are the integers

OpenStudy (anonymous):

...So \(k(k+2) = 323\) for some odd integer \(k\).

OpenStudy (kropot72):

Let x be the lesser of the two numbers. x(x + 2) = 323 \[x ^{2}+2x-323=0\] Find the roots of the above equation to give the positive value for x. Don't forget are two positive numbers and two negative numbers in the solution.

OpenStudy (anonymous):

And be sure to omit the even one since you are given that the integers you seek are odd.

OpenStudy (kropot72):

There is not an even solution to the quadratic.

OpenStudy (anonymous):

Um...yes there are. There are two, and they both happen to be odd. One positive, and the other negative.